OpenStudy (anonymous):

5=12 over x (whats x)

OpenStudy (anonymous):

5 = 12/x multiply both sides by x 5x = 12 divide both sides by 5 x = 12/5

OpenStudy (aravindg):

12/5

OpenStudy (anonymous):

5x = 12 x = 2.4

OpenStudy (anonymous):

x=12/5

OpenStudy (anonymous):

\[5 = \frac{12}{x}\] \[x = \frac{12}{5}\]

OpenStudy (anonymous):

x=12/5